# Spatial vertical regression for spatial panel data: Evaluating the effect of the Florentine tramway’s first line on commercial vitality

### v2

@@ 1 Introduction
In causal inference problems, researchers often deal with panel data with units observed over time. In panel data settings where a single unit or few units are exposed to the treatment starting from a common time point, treatment effects on the treated are often evaluated using the Synthetic Control (SC) method (Abadie et al., 2010; Abadie, 2021; Abadie and Vives-i Bastida, 2022). With this method, the impact of an intervention is evaluated as the difference between the observed value of some outcome and its counterfactual value, imputed using a weighted average of the control units, termed the synthetic control. The popularity of SC methods is growing rapidly, and its field of application ranges from the social sciences, to ecological studies, to policy evaluation. Athey and Imbens (2017) refer to it as “arguably the most important innovation in the policy evaluation literature in the last 15 years” and nowadays (June 2025), the reference paper, Abadie et al. (2010), counts over 7,800 citations on Google Scholar.
@@
In several policy evaluation studies, the treatment is applied at some locations, referred to as treatment sites, and may affect the outcome up to an unknown distance. For instance, as in our case study of the Florentine tramway, the construction of new infrastructure at specific locations in a city may affect the socio-economic development of surrounding areas, at least up to some distance. Currently, there is no guidance on how to deploy SC methods accounting for the spatial propagation of the effects when the treatment is applied to a given spatial location. In practice, one can apply SC methods separately to areas of different size around the treatment site, essentially neglecting the spatial correlation across these areas, which in turn may lead to less precise inferential results. However, exploiting the common spatial patterns across units may allow for information sharing in the imputation of counterfactuals for treated units. This is particularly relevant in contexts with limited information before the intervention occurs, where SC methods are commonly deployed.
@@
In this work, we aim to fill this gap by bridging spatial statistics and synthetic control methodology. We focus on evaluating the effect of a treatment applied on a set of locations on a sequence of treated areas, which are defined as the geographical units of varying distances from the treatment sites. Under the Rubin causal model (Rubin, 1974), we propose a Bayesian approach for imputing the missing potential outcomes that accounts for the presence of spatial correlation among the treated units. We refer to this method as Spatial Vertical Regression or SVR. Adopting from SC methods, we express the missing control potential outcome for a treated unit as a linear combination of the observed outcomes for the control units, up to residual error terms. The key insight is that the relationship of a control unit with the treated areas is expected to be spatially coherent, in that the weight of a control unit in imputing a treated area’s missing potential outcome should be similar across contiguous treated areas. We employ a Gaussian process prior on the coefficients of a control unit, with a spatial correlation matrix defined based on the relative distance of the treated units from the treatment site. Therefore, this specification expresses that the weights given to each control unit in imputing the missing control potential outcome for the treated units may be similar for treated units of similar distance from the treatment locations. Positioned within the Bayesian paradigm, SVR allows us to easily quantify the uncertainty of the estimates, a still debated issue when SC methods are used. From a substantive perspective, the proposed approach is able to provide information on the spread of the treatment effects over space, which may be crucial to policy makers.
@@
We evaluate SVR’s performance through an extensive simulation study. We compare it against a number of alternative approaches, including the original SC and extensions, applied on each treated area separately. The simulation results suggest that the proposed approach performs well in several challenging scenarios. It often leads to a gain in terms of efficiency in comparison to the competitors which ignore the spatial structure, and performs similarly in its absence. It is worth noting that, by leveraging the spatial structure of the data, SVR performs well and exhibits good inference even in scenarios with few pre-treatment time periods, when existing methods like SC or vertical regression without regularization may work poorly or even be infeasible.
@@
Within our motivating context, we evaluate the effect that the construction and operational phases of the Florentine tramway network’s first line had on the surrounding neighborhood’s commercial vitality. This application is relevant as the effect of an urban transport facility may spread in the surrounding areas (Guerra et al., 2012). We study the effect of the tramway for the served areas as well as competition effects with farther areas in the same neighbourhood. We find that the construction and operation of the Florentine tramway network’s first line had small and mostly statistically insignificant effects on the neighborhood’s commercial vitality. More specifically, we observe an increase in the number of stores close to the tramway stops during the construction phase, likely due to the anticipation of the tramway’s operational phase. This effect is very localized, within a two-minute walk from the stops, and diminishes quickly at larger distances. Negative impacts due to the traffic congestion caused by construction work are not present. All the surrounding areas experience a steep decline in the inauguration year of the tramway, followed by a large recovery that granted an overall increase at the end of the observation period.
@@
The paper will follow this outline: Section 2 briefly reviews the literature on causal inference in panel data settings, Section 3 describes the motivating application, Section 4 introduces the causal estimands, and Section 5 introduces the proposed approach. Section 6 presents the simulations studies while Section 7 presents our application results. Finally Section 8 concludes and discusses future research.
@@ 2 Related work
In the last years, a growing number of studies has contributed to the literature on causal inference with panel data, applying and extending the SC method introduced initially by Abadie and Gardeazabal (2003) (see Abadie et al., 2010, 2015, for methodological details) and developing many theoretical advances. See Abadie (2021); Arkhangelsky and Imbens (2023) for extended reviews.
@@
The original SC method imputes the missing control potential outcomes for the treated units in the post-treatment period as convex linear combinations of the control units. The SC method has been characterized as a constrained least squares estimator of a vertical regression model where treated units’ outcomes in each pre-treatment time point are regressed on control units’ outcomes at the same time point (Doudchenko and Imbens, 2016). A number of modifications have been suggested to the original SC method, including SC estimators that relax the convexity restrictions and add intercepts in the vertical regression model (Hsiao et al., 2012; Doudchenko and Imbens, 2016; Ferman and Pinto, 2021), estimators that combine SC methods with other methods, such as matching (Abadie and L’hour, 2021; Kellogg et al., 2021), or Difference-in-Differences (DID) methods (Arkhangelsky et al., 2021), and the augmented SC methods where the SC estimators are augmented using a linear regression of the treated periods outcomes on the pre-treatment outcomes for the control units (Ben-Michael et al., 2021, 2022). A strand of the literature has opted for factor-model approaches that motivate synthetic-control type algorithms and matrix completion methods (Xu, 2017; Amjad et al., 2018; Agarwal et al., 2020; Athey and Imbens, 2021) and approaches with nonlinear panel data models such as the changes-in-changes model introduced by Athey and Imbens (2006) and the distributional SC method (Gunsilius, 2023).
@@
A critical issue in causal inference with panel data, especially in SC settings, concerns the quantification of the uncertainty; there is no consensus regarding the best way to assess the uncertainty of the estimates. A classic sampling-based perspective, where units are viewed as a sample from a larger population, is controversial when the study includes all the units of the population of interest, as it is quite common in SC settings. A design-based or model-based perspective are alternative approaches. A design-based perspective attributes randomness to the treatment assignment mechanism, viewing the potential outcomes as fixed (e.g., Bottmer et al., 2021; Athey and Imbens, 2021). As Bottmer et al. (2021) noted, the design-based perspective underlies the placebo tests that are commonly used in synthetic controls analyses (e.g., Abadie et al., 2010), which requires exchangeability either over units or over time. Conformal inference procedures have been also proposed (e.g., Chernozhukov et al., 2021; Cattaneo et al., 2021), which rely on exchangeability of the residuals from some model over time, and thus, are closely related to the placebo test methods. The model-based perspective views the potential outcomes as random variables. Under the model-based perspective, Bayesian approaches are particularly appealing providing a natural framework to handle the large number of structurally missing potential outcomes with automatic uncertainty quantification, not relying on large sample approximations. Model-based Bayesian inference has been widely used for causal inference in panel data settings, in general, and SC settings, in particular (e.g., Tuomaala, 2019; Kim et al., 2020; Samartsidis et al., 2020; Menchetti and Bojinov, 2020; Pang et al., 2022; Goh and Yu, 2022; Giudice et al., 2022; Ben-Michael et al., 2023; Fernández-Morales et al., 2025). See also Martinez and Vives-i Bastida (2022) for a discussion on Bayesian and Frequentist inference for SCs.
@@
In this paper, we use a model-based Bayesian approach to inference, which allows us to naturally incorporate prior substantive knowledge about the spatial data structure into the causal analysis. The key feature of the SVR approach we propose is the use of a Gaussian process prior distribution to borrow information across spatially correlated units in a panel data setting. SVR has similarities with the methods proposed by Vega and Nethery (2023) and Ben-Michael et al. (2023), both of which employ latent factor models for the potential outcomes. In a similar vein to us, the authors of these manuscripts deal with the spatial structure of the data using prior distributions in a Bayesian framework. Specifically, Vega and Nethery (2023) specify spatially-informed priors on the unit-specific factor loadings, and Ben-Michael et al. (2023) use multitask Gaussian Processes, a non-parametric Bayesian approach, to generalize low-rank factor models with a prior that encourages smoothness in the underlying factors. In contrast, our work is set within the SC framework which specifies that the control potential outcomes for the treated units are a unit-specific linear combination of the control potential outcomes for the control units. We capture spatial correlation across the treated units by employing a Gaussian process prior distribution on the coefficients of each control unit in the vertical regression model for the treated areas. On a different vein, Fernández-Morales et al. (2025) use Bayesian shrinkage priors to penalize the inclusion of short-distance control units to avoid biased estimates due to spillover effects.
@@
Our work is also related to the literature on causal inference with spatially correlated data in panel data settings, which often arise in environmental studies, possibly generating spillover effects (e.g., Delgado and Florax, 2015; Sanford, 2018; Bardaka et al., 2018; Antonelli and Beck, 2020; Butts, 2021; Hettinger et al., 2023) and the literature on SC methods with multiple outcomes (see, e.g., Billmeier and Nannicini, 2013; Acemoglu et al., 2016; Cunningham and Shah, 2018; Kasy and Lehner, 2023; Ben-Michael et al., 2023; Sun et al., 2023; Tian et al., 2023). In causal inference with spatial panel data, focus is usually on accounting for spatial correlation or spatial spillover effects, whereas we are interested in assessing how the treatment effects propagate through space. In the presence of multiple outcomes, SC methods are often applied separately to each outcome variable. Recently, Ben-Michael et al. (2023); Sun et al. (2023); Tian et al. (2023) proposed methods for estimating effects on multiple outcomes simultaneously. These methods are based on low-rank factor models with shared latent factors across different outcomes, and it is not completely clear which aspect of the data informs the correlation among outcomes. In the presence of spatial outcomes, we exploit the external information about the spatial structure in a principled way.
@@ 3 Florence’s first tramway line
The first line of the Florentine transit network began an ambitious urban renovation, improving accessibility in one of Italy’s most traffic-congested cities. Although plans for an integrated tramway date back to the 1990s, construction began in 2006, culminating in an inaugural run in 2010. As of 2024, two lines are operational, and three more are in development, slated for completion by 2030.
@@
Investments in transportation infrastructure, especially in light rail systems, spark debate in urban and transport economics (e.g., Cervero and Landis, 1993, Landis et al., 1995, Baum-Snow and Kahn, 2000, Hess and Almeida, 2007, Pan, 2013, Papa and Bertolini, 2015). Such projects aim to reduce travel time and congestion while often revitalizing public spaces and commercial areas. These benefits can be more pronounced where there is a well-connected transit network, as shown by Mejia-Dorantes et al. (2012) and Credit (2018). Regarding economic outcomes, research commonly focuses on the positive correlation between improved transport and real estate prices (Pagliara and Papa, 2011; Yan et al., 2012). Some studies also consider commercial vitality: Budiakivska and Casolaro (2018) find positive effects on property values after Florence’s first tramway line, while Grossi et al. (2025) examine direct and spillover effects on local commerce.
@@
Our goal is to evaluate the impact of the tramway on commercial vitality during the construction and operational phases of the tramway line, and, importantly, how it diffuses spatially. In an evaluation study, it is crucial to consider both the construction and operational phases, since reduced accessibility during construction may displace businesses, especially those near the site. Nevertheless, once the tramway becomes operational, nearby areas may enjoy better accessibility and surroundings (Credit, 2018; Pogonyi et al., 2021).
@@
Figure 1 shows the study area: the black line and points indicate the first Florentine tramway line and stops, respectively. As we aim to assess how the treatment effect spreads, we define a series of buffers for the tramway line. These buffers mark catchment areas within 600 seconds walking distance from any tram stop, in two-minute increments, resulting in five buffers per line representing the five treated areas. Walking distances are calculated using the API for isochrones from Openroute service and the osrm package in R (Blanchet and Lovelace, 2023). We draw 11 hypothetical, alternative tramway lines and corresponding stops, shown in the figure in white color. These lines and stops will be the basis for acquiring a synthetic control. They are chosen either because they follow lines that have been constructed after our study period, or similar paths to the realized one (see also Grossi et al., 2025, for more description on these alternative tramway lines). We define control areas in a similar manner to the treated areas.
@@
Figure 1: Catchment areas around the tramway stops, buffers represent areas within 120, 240, 360, 480 and 600 seconds of walking from a tramway stop, shown in red to yellow color. Black points and line represent the realized stops and tramway line, respectively, while white points and lines represent the hypothetical stops and tramway lines.
@@
Our primary outcome to capture commercial vitality is the number of stores within each catchment area. Store locations come from ISTAT’s Statistical Archive of Active Firms (SAAF) at a yearly level for 1996–2016. Construction of the tramway began in early 2006 and ended in 2009. Therefore, the pre-treatment period includes only 10 time periods. The outcome of the five treated areas is defined as the number of stores within the corresponding buffers at the five different distances from the tram stops. For the control areas, the outcome is defined as the average number of stores in the five buffers around the hypothetical tramway line. Figure 2 presents time series for treated and control units, showing limited variation in store counts and a common citywide trend, largely driven by macroeconomic factors. Vertical lines represent the beginning of the construction and operational phases of the tramway line.
@@
Figure 2: Number of Stores in Treated and Control Areas Across Time. Treated units, shown in orange, correspond to five catchment areas within 120, 240, 360, 480 and 600 seconds of walking distance from any of the realized tramway stops. Control units, shown in light blue, correspond to average number of stores in their catchment areas, corresponding to the 10 hypothetical tramway lines.
4 Causal Framework: towards spatial panel models

### v3

@@ 1 Introduction
In causal inference, researchers often deal with panel data where one or a few units experience the treatment starting from a common time point, while the remaining units remain control throughout. Treatment effects on the treated units are often evaluated using the Synthetic Control (SC) method (Abadie et al., 2010; Abadie, 2021; Abadie and Vives-i Bastida, 2022). With this method, the impact of an intervention is evaluated as the difference between the observed value of the outcome and its counterfactual value, imputed using a weighted average of outcomes of the control units, termed the synthetic control. The popularity of SC methods is growing rapidly, and its field of application ranges from the social sciences, to ecological studies, to policy evaluation. Athey and Imbens (2017) refer to it as “arguably the most important innovation in the policy evaluation literature in the last 15 years” and nowadays (March 2026), the reference paper, Abadie et al. (2010), counts over 8,600 citations on Google Scholar.
@@
In several policy evaluation studies, the treatment is applied at some locations, referred to as treatment sites, and may affect the outcome up to an unknown distance. For instance, as in our case study of the Florentine tramway, the construction of new infrastructure at specific locations in a city may affect the socio-economic development of surrounding areas up to some distance. Currently, there is no guidance on how to deploy SC methods accounting for the spatial propagation of the effects when the treatment is applied to a given spatial location. In practice, one can apply SC methods separately to areas of different size around the treatment site, essentially neglecting the spatial correlation across these areas, which in turn may lead to less precise inferential results. However, exploiting the common spatial patterns across units may allow for information sharing in the imputation of counterfactuals for treated units. This is particularly relevant in contexts with few time periods with observed data before the treatment occurs, where SC methods are commonly deployed.
@@
In this work, we aim to fill this gap by bridging spatial statistics and synthetic control methodology. We focus on evaluating the effect of a treatment applied on a set of locations on a sequence of treated areas, which are defined as the geographical units of varying distances from the treatment sites. Under the Rubin causal model (Rubin, 1974), we propose a Bayesian approach for imputing the missing potential outcomes that accounts for the presence of spatial correlation among the treated units. We refer to this method as Spatial Vertical Regression or SVR. The term vertical regression was first introduced by Athey et al. (2021) to describe synthetic control-based methods. Adopting this approach, we express the missing control potential outcome for a treated unit as a linear combination of the observed outcomes for the control units, up to residual error terms. The key insight is that the relationship of a control unit with the treated areas is expected to be spatially coherent, in that the weight of a control unit in imputing a treated area’s missing potential outcome should be similar across contiguous treated areas. SVR encodes this by employing a Gaussian process prior on the coefficients of each control unit with a spatial correlation matrix. This specification expresses that the weights given to each control unit in imputing the missing control potential outcome for the treated units may be similar for treated units of similar distance from the treatment locations. Positioned within the Bayesian paradigm, SVR allows us to easily quantify the uncertainty of the estimates, a still debated issue when SC methods are used. From a substantive perspective, the proposed approach is able to provide information on the spread of the treatment effects over space, which may be crucial to policy makers.
@@
We evaluate SVR’s performance and compare it against alternative approaches through an extensive simulation study. The simulation results suggest that SVR leads to efficiency gains in comparison to the competitors when spatial structure is present, and performs similarly in its absence. By leveraging the spatial structure of the data, SVR performs well and exhibits good inferential properties even in scenarios with few pre-treatment time periods, when existing methods may work poorly or even be infeasible.
@@
Our motivating context is the evaluation of the effect that the construction and operational phases of the Florentine tramway network’s first line had on the surrounding neighborhood’s commercial vitality. This application is relevant as the effect of an urban transport facility may spread in the surrounding areas (Guerra et al., 2012). We study the effect of the tramway for the served neighbourhood at different distances from the stops. We find that the construction and operation of the tramway line had small and mostly statistically insignificant effects on the neighborhood’s commercial vitality. More specifically, we observe an increase in the number of stores close to the tramway stops during the construction phase, likely due to the anticipation of the tramway’s operational phase. This effect is very localized, within a two-minute walk from the stops, and diminishes quickly at larger distances. Negative impacts due to traffic congestion caused by construction work are not present. Areas at farther distances experience a steep decline in the inauguration year of the tramway, followed by a large recovery that granted an overall increase at the end of the observation period.
@@
The paper will follow this outline: Section 2 briefly reviews the literature on causal inference with panel data, Section 3 describes the motivating application, Section 4 introduces the causal estimands, and Section 5 introduces the proposed approach. Section 6 presents the simulations studies while Section 7 presents our application results. Finally Section 8 concludes and discusses future research.
@@ 2 Related work
In the last years, a growing number of studies has contributed to the literature on causal inference with panel data, where a number of units are followed over time, and a subset of them is exposed to a treatment starting at a specific time point. Recent work applies, extends, or provides theoretical advances to the SC method introduced initially by Abadie and Gardeazabal (2003) (see Abadie et al., 2010, 2015, for methodological details). See Abadie (2021) and Arkhangelsky and Imbens (2023) for extended reviews.
In panel data settings, estimating causal effects often relies on estimating counterfactual control outcomes for the treated units in the post-intervention period. Vertical Regression (VR) methods are often employed towards this goal. The term “vertical regression” was first introduced by Athey and Imbens (2021) to describe methods in which treated units’ potential outcomes under control are expressed as a linear combination of the outcomes of control units. The coefficients of this linear combination are estimated by regressing treated units’ outcomes on control units’ outcomes at each pre-treatment time point. These estimated coefficients are then applied to the control units’ outcomes in the post-treatment period to predict the treated units’ missing potential outcomes under control. The original SC method imputes the missing control potential outcomes for the treated units in the post-treatment period as convex linear combinations of the control units. Therefore, the SC method has been characterized as a constrained least squares estimator of a vertical regression model.
In turn, vertical regression methods can be viewed as a modification of the SC method that relax the convexity restrictions and add intercepts (Hsiao et al., 2012; Doudchenko and Imbens, 2016; Ferman and Pinto, 2021).
@@
A number of additional modifications have been suggested to the original SC method, including estimators that combine SC methods with matching (Abadie and L’hour, 2021; Kellogg et al., 2021) or difference-in-differences (Arkhangelsky et al., 2021), and the augmented SC estimators which also use a linear regression of the treated periods outcomes on the pre-treatment outcomes for the control units (Ben-Michael et al., 2021, 2022). A strand of the literature has opted for factor-model approaches that motivate synthetic-control type algorithms and matrix completion methods (Xu, 2017; Amjad et al., 2018; Agarwal et al., 2020; Athey and Imbens, 2021) and approaches with nonlinear panel data models such as the changes-in-changes model introduced by Athey and Imbens (2006) and the distributional SC method (Gunsilius, 2023).
@@
In synthetic control (SC) settings, quantifying uncertainty around the counterfactual trajectory is non-trivial, as standard errors are not available off-the-shelf. A classic sampling-based perspective, where units are viewed as a sample from a larger population, is controversial when the study includes all the units of the population of interest, as it is quite common in SC settings. A design-based or model-based perspective are alternative approaches. A design-based perspective attributes randomness to the treatment assignment mechanism, viewing the potential outcomes as fixed (e.g., Bottmer et al., 2024; Athey and Imbens, 2021). As Bottmer et al. (2024) noted, the design-based perspective underlies the placebo tests that are commonly used in synthetic controls analyses (e.g., Abadie et al., 2010), which requires exchangeability either over units or over time. Conformal inference procedures have also been proposed (e.g., Chernozhukov et al., 2021; Cattaneo et al., 2021), which rely on exchangeability of the residuals from some model over time, and thus, are closely related to the placebo test methods. The model-based perspective views the potential outcomes as random variables. Under this perspective, Bayesian approaches provide a natural framework to handle the large number of structurally missing potential outcomes with automatic uncertainty quantification, without relying on large sample approximations. Model-based Bayesian inference has been widely used for causal inference with panel data in general, and in SC settings in particular (e.g., Tuomaala, 2019; Kim et al., 2020; Samartsidis et al., 2020; Menchetti and Bojinov, 2020; Pang et al., 2022; Goh and Yu, 2022; Giudice et al., 2022; Ben-Michael et al., 2023; Fernández-Morales et al., 2025). See also Martinez and Vives-i Bastida (2022) for a discussion on Bayesian and Frequentist inference for SCs.
@@
In this paper, we use a model-based Bayesian approach to inference, which allows us to naturally incorporate prior substantive knowledge about the spatial data structure into the causal analysis. The key feature of the SVR approach we propose is the use of a Gaussian process prior distribution to borrow information across spatially adjacent units in a panel data setting. SVR has similarities with the methods proposed by Vega and Nethery (2023) and Ben-Michael et al. (2023), both of which employ latent factor models for the potential outcomes. In a similar vein to us, the authors of these manuscripts deal with the spatial structure of the data using prior distributions in a Bayesian framework. Specifically, Vega and Nethery (2023) specify spatially-informed priors on the unit-specific factor loadings, and Ben-Michael et al. (2023) use multitask Gaussian Processes, a non-parametric Bayesian approach, to generalize low-rank factor models with a prior that encourages smoothness in the underlying factors. In contrast, our work is set within the SC framework which specifies that the control potential outcomes for the treated units are a unit-specific linear combination of the control potential outcomes for the control units. We capture spatial correlation across the treated units by employing a Gaussian process prior distribution on the coefficients of each control unit in the vertical regression model for the treated areas. On a different vein, Fernández-Morales et al. (2025) use Bayesian shrinkage priors to penalize the inclusion of short-distance control units to avoid biased estimates due to spillover effects.
@@
Our work is also related to the literature on causal inference with spatially correlated data in panel data settings, which often arise in environmental studies, possibly generating spillover effects (e.g., Delgado and Florax, 2015; Sanford, 2018; Bardaka et al., 2018; Antonelli and Beck, 2020; Butts, 2021; Hettinger et al., 2023) and the literature on SC methods with multiple outcomes (see, e.g., Billmeier and Nannicini, 2013; Acemoglu et al., 2016; Cunningham and Shah, 2018; Kasy and Lehner, 2023; Ben-Michael et al., 2023; Sun et al., 2023; Tian et al., 2023). In causal inference with spatial panel data, focus is usually on accounting for spatial correlation or spatial spillover effects, whereas we are interested in assessing how the treatment effects propagate across space. In the presence of multiple outcomes, SC methods are often applied separately to each outcome variable. Recently, Ben-Michael et al. (2023); Sun et al. (2023); Tian et al. (2023) proposed methods for estimating effects on multiple outcomes simultaneously. These methods are based on low-rank factor models with shared latent factors across different outcomes, and it is not completely clear which aspect of the data informs the correlation among outcomes. In the presence of spatial outcomes, we exploit the external information about the spatial structure in a principled way.
@@ 3 Florence’s first tramway line
The first line of the Florentine transit network began an ambitious urban renovation, improving accessibility in one of Italy’s most traffic-congested cities. Although plans for an integrated tramway date back to the 1990s, construction began in 2006, culminating in an inaugural run in 2010. As of 2026, two lines are operational, and three more are in development, slated for completion by 2030.
@@
Investments in transportation infrastructure, especially in light rail systems, spark debate in urban and transport economics (e.g., Cervero and Landis, 1993, Landis et al., 1995, Baum-Snow and Kahn, 2000, Hess and Almeida, 2007, Pan, 2013, Papa and Bertolini, 2015). Such projects aim to reduce travel time and congestion while often revitalizing public spaces and commercial areas. These benefits can be more pronounced where there is a well-connected transit network, as shown by Mejia-Dorantes et al. (2012) and Credit (2018).
Research commonly focuses on the positive correlation between improved transport and real estate prices (Pagliara and Papa, 2011; Yan et al., 2012). Complementary to that literature, some studies focus on the effects of changes in transportation systems on commercial vitality and location choices (Ray, 2015; Nilsson and Smirnov, 2016; Neto et al., 2025). Some work focuses specifically on the causal effects of the construction of the first line of the Florentine tramway. Grossi et al. (2025) study its effect on the commercial environment whereas Budiakivska and Casolaro (2018) find an overall positive effect on real estate prices, coherent with the literature.
@@
We contribute to this literature by evaluating the impact of the tramway on commercial vitality during the construction and operational phases of the tramway line, and, importantly, how it diffuses spatially at different distances from the tram stops. In an evaluation study, it is crucial to consider both the construction and operational phases, since reduced accessibility during construction may displace businesses, especially those near the site. Nevertheless, once the tramway becomes operational, nearby areas may enjoy better accessibility and surroundings (Credit, 2018; Pogonyi et al., 2021).
@@
Figure 1 shows the study area: the black line and points indicate the first Florentine tramway line and stops, respectively. As we aim to assess how the treatment effect spreads across space, we define five catchment areas within 600 seconds walking distance from any tram stop, in two-minute increments. Walking distances are calculated using the API for isochrones from Openroute service and the osrm package in R (Blanchet and Lovelace, 2023). We draw 11 hypothetical, alternative tramway lines and corresponding stops, shown in the figure in white color. These lines and stops will be the basis for acquiring a synthetic control. They are chosen either because they follow lines that have been constructed after our study period, or similar paths to the realized one (see also Grossi et al., 2025, for more description on these alternative tramway lines). We define control catchment areas in a similar manner to the treated ones.
@@
Figure 1: Catchment areas around the tramway stops representing areas within 120, 240, 360, 480 and 600 seconds of walking from a tramway stop, shown in red to yellow color. Black points and line represent the realized stops and tramway line, respectively, while white points and lines represent the hypothetical stops and tramway lines.
@@
Our primary outcome to capture commercial vitality is the number of stores within each catchment area. Store locations come from yearly data from ISTAT’s Statistical Archive of Active Firms (SAAF) for 1996–2016. Construction of the tramway began in early 2006 and ended in 2009. Therefore, the pre-treatment period includes only 10 time periods. The outcome of the five treated areas is defined as the number of stores within the corresponding catchment areas at the five different distances from the tram stops. For the control areas, the outcome is defined as the average number of stores in the five catchment areas around the hypothetical tramway line. Figure 2 presents time series for treated and control units, showing limited variation in store counts and a common citywide trend, largely driven by macroeconomic factors. Vertical lines represent the beginning of the construction and operational phases of the tramway line.
@@
Figure 2: Number of Stores in Treated and Control Areas Across Time. Treated units, shown in orange, correspond to five catchment areas within 120, 240, 360, 480 and 600 seconds of walking distance from any of the realized tramway stops. Control units, shown in light blue, correspond to average number of stores in their catchment areas, corresponding to the $1$ hypothetical tramway lines.
4 Causal framework on spatial panel data